# On the comparison of ultradifferentiable weight function settings

### v1

@@ 2.3. Relevant growth relations between weight functions
    (2.3) $\exists\;t_{0}>0\;\exists\;\epsilon>0\;\forall\;t\geq t_{0}:\;\;\;\sigma(t)\geq\epsilon,$
@@
    (2.3) is a mild requirement; e.g. it holds if $\sigma\neq 0$ is assumed to be non-decreasing. Moreover it holds if $\lim_{t\rightarrow+\infty}\sigma(t)=+\infty$ and this has been a basic and natural growth requirement when studying weight functions even in a very general setting in [12].
@@
All properties listed in Section 2.2 except $(\omega_{\operatorname{sub}})$ and the convexity property $(\omega_{4})$ are automatically preserved under $\sim$; see also [29, Sect. 3.2]. Indeed, the effects on $(\omega_{\operatorname{sub}})$ when passing to an equivalent weight and the relation to $(\alpha_{0})$ are investigated and revisited in Section 2.5 whereas the failure of $(\omega_{4})$ is studied in detail in Section 7 where a (counter-)example is constructed.
@@ 2.4. On the comparison of crucial growth properties
    $(\alpha_{0})$ implies $(\omega_{1})$ (choose $\lambda=2$).
@@
    Also $(\omega_{\operatorname{sub}})$ implies $(\omega_{1})$ (take $x=y$) and with $x=y=0$ we get $\omega(0)\leq 2\omega(0)$ which yields $\omega(0)\geq 0$.
@@
    In the literature sometimes the following condition is used (see e.g. [7, Lemma 1.2]):
@@
    Set $s=t$ and so get immediately $(\omega_{1})$ when choosing the constant $2L$. If $\omega$ is non-decreasing, then (2.5) is equivalent to $(\omega_{1})$: Indeed, if $\omega$ has $(\omega_{1})$ then $\omega(s+t)\leq\omega(2\max\{s,t\})\leq L\omega(\max\{s,t\})+L\leq L(\omega(s)+\omega(t))+L$ and so (2.5) holds with the same $L$.
@@
    Let $\omega$ be non-decreasing (and such that $\omega\neq 0$). Then concavity on $0,+\infty)$ and normalization can never be satisfied at the same time.
    The same technical problem appears for $(\alpha_{0})$, see also [30, Sect. 4.1] and the references therein: Assume that there exists $t_{0}>0$ such that $\omega(t)=0$ for all $t\in[0,t_{0}]$, then $C\lambda\omega(t)=0$ for all $C,\lambda\geq 1$ and all $0<t\leq t_{0}$. But, since $\omega\neq 0$ and $\omega$ is non-decreasing, we have $\omega(\lambda t)>0$ as $\lambda\rightarrow+\infty$ for any $t>0$.
    And similarly also normalization and $(\omega_{\operatorname{sub}})$ can never be achieved at the same time: If there exists $t_{0}>0$ such that $\omega(t)=0$ for all $t\in[0,t_{0}]$ then for all $0<s\leq t_{0}$ and all $t\geq 0$ we get $\omega(t)\leq\omega(s+t)\leq\omega(s)+\omega(t)=\omega(t)$. But this implies $\omega=0$.
@@
      $\forall\;a>1\;\exists\;L\geq 1\;\forall\;t\geq 0:\;\;\;\omega(at)\leq L\omega(t)+L$
@@
Proof. The particular implications $(i)\Rightarrow(ii)$ and $(iii)\Rightarrow(i)$ are clear (set $a:=2$) and $(ii)\Rightarrow(iii)$ follows by iteration; concerning $(\omega_{6})$ see also [28, Rem. 2.2]. However, the assumption $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$ is not required necessarily for these equivalences. ∎
@@ 2.5. On subadditivity-like conditions
We recall the following equivalent reformulations of $(\alpha_{0})$; see [21, Lemma 1], [12, Prop. 2.23] and finally [30, Lemma 4.1]:
@@
    Assumption $\omega(0)\geq 0$ in [30, Lemma 4.1] is superfluous.
    To be formally correct, in the proof of [30, Lemma 4.1 $(ii)\Rightarrow(iii)$] one shall assume $s>0$ and note that the second part of [12, $(12)$] is trivial when $s=0$.
    In [30] equivalence is defined via O-growth relations like in (2.2), see [30, Sect. 2.4, p. 406], and which coincides with the above given notion $\sim$ since (2.3) holds by $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$ as a basic requirement in this work.
    However, in [30, Lemma 4.1] it was only assumed that $\omega$ is non-decreasing but this is sufficient to get (2.3) when $\omega\neq 0$ and as seen before both notions of equivalence coincide in this case.
@@
$(\alpha_{0})$ should be compared with $(\alpha_{1})$ in [23]. Note that in [23, Prop. 1.1] it has been claimed that $(\alpha_{1})$ is equivalent to $(\alpha_{0})$ but which fails in general and $(\alpha_{1})$ is weaker. However, the difference between both requirements is subtle, we refer to [30] for more comments and explanations.
@@
For arbitrary $\omega:[0,+\infty)\rightarrow[0,+\infty)$ introduce $\omega^{\iota}:(0+\infty)\rightarrow[0,+\infty)$ given by $\omega^{\iota}(t):=\omega(1/t)$. So $\omega^{\iota}$ is non-increasing if and only if $\omega$ is non-decreasing, $\lim_{t\rightarrow 0}\omega^{\iota}(t)=\lim_{t\rightarrow+\infty}\omega(t)$ and $\lim_{t\rightarrow+\infty}\omega^{\iota}(t)=\lim_{t\rightarrow 0}\omega(t)$ if the limits exist.
@@
Let $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing. If $\omega$ satisfies (2.6), then
@@
and then apply [12, Prop. 2.24]. Note that (2.7) is equivalent to the existence of universal constants $C_{1},C,D\geq 1$ not depending on $\beta$ such that $C_{1}\omega^{\iota}(s)+D\geq\frac{1}{C_{2}}\min\left\{1,\frac{t^{\beta}}{s^{\beta}}\right\}\omega^{\iota}(s)$ holds for all $s,t\in(0,+\infty)$ (set $C:=\max\{D/C_{1},C_{1}C_{2}\}$).
@@
Second, note that (2.6) gives with $t^{\prime}:=\frac{1}{\lambda t}$:
$\exists\;C,D\geq 1\;\forall\;\lambda\geq 1\;\forall\;t^{\prime}>0:\;\;\;\omega^{\iota}(t^{\prime})=\omega(1/t^{\prime})\leq C\lambda\omega(1/(\lambda t^{\prime}))+D\lambda=C\lambda\omega^{\iota}(\lambda t^{\prime})+D\lambda.$
Let now $0<t<s$ be arbitrary and choose $\lambda=\frac{s}{t}>1$, then the above estimate gives
$D\lambda+C\lambda\omega^{\iota}(s)=D\lambda+C\lambda\omega^{\iota}(\lambda t)\geq\omega^{\iota}(t),$
hence for any $\beta\geq 1$:
$D+C\omega^{\iota}(s)\geq\frac{t}{s}\omega^{\iota}(t)\geq\frac{t^{\beta}}{s^{\beta}}\omega^{\iota}(t)=\min\left\{1,\frac{t^{\beta}}{s^{\beta}}\right\}\omega^{\iota}(t).$
@@
And if there exists some $t_{0}>0$ such that $\omega(t)=0$ for all $t\in[0,t_{0}]$, then $\omega^{\iota}(t)=0$ for all $t\geq t_{0}^{-1}$. Applying the previous estimate to $s:=t_{0}^{-1}$ gives $\omega^{\iota}(\lambda_{1}t+\lambda_{2}t_{0})\leq A^{2}\lambda_{1}\omega^{\iota}(t)+A^{2}$ for all $t\in(0,t_{0}]$ and $\lambda_{1},\lambda_{2}\geq 0$ satisfying $\lambda_{1}+\lambda_{2}=1$.
Remark that the proof and the conclusion from [29, Lemma 3.8.1 $(2)$] fails since super-additivity for $\omega^{\iota}$ would imply $2\omega^{\iota}(t)\leq\omega^{\iota}(2t)\leq\omega^{\iota}(t)$ for all $t\in(0,+\infty)$, a contradiction.
@@ 2.6. Comment on weights being defined on $\mathbb{R}^{d}$ or $\mathbb{C}^{d}$
2.7. The growth index $\gamma(\omega)$
We briefly recall the definition of the growth index $\gamma(\omega)$, see [12, Sect. 2.3] and the references therein. Let in this section $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing and such that $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$; i.e. $\omega$ is a weight function as in [12, Sect. 2.3]. For $\gamma>0$ we say that $\omega$ has property $(P_{\omega,\gamma})$ if
@@ 2.7. The growth index $\gamma(\omega)$
(2.8) $\gamma(\omega):=\sup\{\gamma>0:(P_{\omega,\gamma})\;\;\text{is satisfied}\},$
@@
Let $\omega$ be as before. We recall some connections between $\gamma(\omega)$ and sub-additivity-like conditions for $\omega$, see also [30, Sect. 6]:
@@
    $\kappa_{\omega}$ is concave, continuous and we have $\kappa_{\omega}(0)=\omega(0)\geq 0$ (i.e. $\kappa_{\omega}(0)=0$ if and only if $\omega(0)=0$). Hence it is known that also $(\omega_{\operatorname{sub}})$ holds for $\kappa_{\omega}$; see e.g. [29, Lemma 3.8.1 $(1)$].
@@
2.8. Different notions of weight functions
In this section we revisit the different notions of weight functions appearing in the literature.
Definition 2.7.
    continuous,
    non-decreasing,
    $\omega(0)=0$, and
    $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$.
Note: Occasionally in [12] we have even worked with a more general notion. Indeed, the assumptions continuity and $\omega(0)=0$ can be removed when working with abstract growth indices; see [12, Sect. 2.2-2.5] and Section 2.7.
Definition 2.8.
Moreover:

### v2

@@ 2.3. Relevant growth relations between weight functions
    (2.2) $\exists\;t_{0}\geq 0\;\exists\;\epsilon>0\;\forall\;t\geq t_{0}:\;\;\;\sigma(t)\geq\epsilon,$
@@
[…]
@@
  • $(v)$
    In view of $(ii)$, for any $\sigma$ satisfying (2.2) we get that $(\omega_{1})$ is equivalent to $\omega(2t)\leq L\omega(t)+L$ for some $L\geq 1$ and all $t\geq 0$ and this estimate is appropriate even if (2.2) fails.
    On the other hand, for $(\omega_{2})$ and $(\omega_{5})$ this technical observation is not required since $t\mapsto t$ is obviously satisfying (2.2).
All properties listed in Section 2.2 except $(\omega_{\operatorname{sub}})$ and the convexity property $(\omega_{4})$ are automatically preserved under $\sim$; see also [28, Sect. 3.2]. Indeed, the effects on $(\omega_{\operatorname{sub}})$ when passing to an equivalent weight and the relation to $(\alpha_{0})$ are investigated and revisited in Section 2.5 whereas the failure of $(\omega_{4})$ is studied in detail in Section 7 where a (counter-)example is constructed.
@@ 2.4. On the comparison of crucial growth properties
    If $\omega$ satisfies (2.2), then $(\alpha_{0})$ implies $(\omega_{1})$ (choose $\lambda=2$).
@@
    If $\omega$ satisfies (2.2), then also $(\omega_{\operatorname{sub}})$ implies $(\omega_{1})$ (take $x=y$) and with $x=y=0$ we get $\omega(0)\leq 2\omega(0)$ which yields $\omega(0)\geq 0$.
@@
    In the literature sometimes instead of $(\omega_{1})$ the following condition is used (see e.g. [7, Lemma 1.2]):
@@
    When (2.2) holds, then set $s=t$ and get immediately $(\omega_{1})$ when choosing the constant $2L$. If $\omega$ is non-decreasing, then (2.4) is equivalent to $(\omega_{1})$: Indeed, if $\omega$ has $(\omega_{1})$ then $\omega(s+t)\leq\omega(2\max\{s,t\})\leq L\omega(\max\{s,t\})+L\leq L(\omega(s)+\omega(t))+L$ and so (2.4) holds with the same $L$.
@@
    Let $\omega$ be non-decreasing (and $\omega\neq 0$) and assume that there exists some $t_{0}>0$ such that $\omega(t)=0$ for all $t\in[0,t_{0}]$. In particular, this last assumption holds when $\omega$ is normalized; i.e. $\omega(t)=0$ for $t\in[0,1]$. Then several growth conditions are violated:
    • $(a)$
      Concavity fails: This condition means $\omega(\lambda t+(1-\lambda)s)\geq\lambda\omega(t)+(1-\lambda)\omega(s)$ for all $s,t\geq 0$ and $\lambda\in[0,1]$. Choose $s\in[0,t_{0})$, $t\in(t_{0},+\infty)$ with $\omega(t)>0$ and such $t$ exists since $\omega$ is non-decreasing and $\omega\neq 0$. Then there exists $\lambda_{0}\in(0,1)$ such that $\lambda_{0}s+(1-\lambda_{0})t=t_{0}$ and thus $0=\omega(t_{0})\geq(1-\lambda_{0})\omega(t)>0$ follows, a contradiction.
    • $(b)$
      $(\alpha_{0})$ fails (see also [29, Sect. 4.1] and the references therein): $C\lambda\omega(t)=0$ holds for all $C,\lambda\geq 1$ and all $0<t\leq t_{0}$ but, since $\omega\neq 0$ and $\omega$ is non-decreasing, we have $\omega(\lambda t)>0$ as $\lambda\rightarrow+\infty$ for any $t>0$ fixed.
    • $(c)$
      $(\omega_{\operatorname{sub}})$ fails: For all $0<s\leq t_{0}$ and all $t\geq 0$ we get $\omega(t)\leq\omega(s+t)\leq\omega(s)+\omega(t)=\omega(t)$ which implies $\omega=0$; again a contradiction.
    Indeed, $(a)$ and $(b)$ are even valid if (2.2) and $\omega(t)=0$ for all $t\in[0,t_{0}]$.
@@
      $\forall\;a>1\;\exists\;L\geq 1\;\forall\;t\geq 0:\;\;\;\omega(at)\leq L\omega(t)+L.$
@@
Proof. The particular implications $(i)\Rightarrow(ii)$ and $(iii)\Rightarrow(i)$ are clear (set $a:=2$) and $(ii)\Rightarrow(iii)$ follows by iteration: Let $b>1$ be arbitrary, then choose $n\in\mathbb{N}_{>0}$ such that $a^{n}\geq b$ with $a$ appearing in $(ii)$, and iterate the corresponding estimate in $(ii)$ $n$-times. (Concerning $(\omega_{6})$ see also [34, Rem. 2.2] but, however, the assumption $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$ made there is not required necessarily for the proof of these equivalences.) ∎
@@ 2.5. On subadditivity-like conditions
We recall the following equivalent reformulations of $(\alpha_{0})$; see [21, Lemma 1], [12, Prop. 2.23] and finally [29, Lemma 4.1]:
@@
    Assumption $\omega(0)\geq 0$ in [29, Lemma 4.1] is superfluous.
    To be formally correct, in the proof of [29, Lemma 4.1 $(ii)\Rightarrow(iii)$] one shall assume $s>0$ and note that the second part of [12, $(12)$] is trivial when $s=0$.
    In [29] equivalence is defined via O-growth relations like in (2.1), see [29, Sect. 2.4, p. 406], and which coincides with the above given notion $\sim$ since (2.2) holds by $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$ as a basic requirement in this work.
    However, in [29, Lemma 4.1] it was only assumed that $\omega$ is non-decreasing but this is sufficient to get (2.2) when $\omega\neq 0$ and as seen before both notions of equivalence coincide in this case.
@@
$(\alpha_{0})$ should be compared with $(\alpha_{1})$ in [23] which reads as follows:
(2.6) $\sup_{\lambda\geq 1}\limsup_{t\rightarrow+\infty}\frac{\omega(\lambda t)}{\lambda\omega(t)}<+\infty.$
Note that in [23, Prop. 1.1] it has been claimed that (2.6) is equivalent to $(\alpha_{0})$ but $(\alpha_{1})$ is weaker and the difference is subtle: $(\alpha_{0})$ requires for the quotient under consideration uniform boundedness for all $t\geq t_{0}$ and for all $\lambda\geq 1$ and $t_{0}$ must not depend on $\lambda$ whereas for $(\alpha_{1})$ it is sufficient to have boundedness for each $\lambda\geq 1$ separately but here the values for $t$ may depend on $\lambda$, i.e. having the estimate only for all $t\geq t_{\lambda}$, cf. [29, $(5.1)$]. We refer to [29] for more comments and explanations and also to the discussion in [28, Sect. 3.8, p. 26-27]: There the (counter-)example given by $\omega(t):=t\log(t)$ for $t>1$ and $\omega(t):=0$ for $t\in[0,1]$ is studied.
@@
For arbitrary $\omega:[0,+\infty)\rightarrow[0,+\infty)$ introduce $\omega^{\iota}:(0+\infty)\rightarrow[0,+\infty)$ given by $\omega^{\iota}(t):=\omega(1/t)$. So $\omega^{\iota}$ is non-increasing if and only if $\omega$ is non-decreasing, $\lim_{t\rightarrow 0}\omega^{\iota}(t)=\lim_{t\rightarrow+\infty}\omega(t)$ and $\lim_{t\rightarrow+\infty}\omega^{\iota}(t)=\lim_{t\rightarrow 0}\omega(t)$ if the limits exist. Moreover, $\omega^{\iota}=0$ if and only if $\omega=0$.
@@
Now we are ready to prove the main result; one of the referees has modified in the report the original proof and was able to replace the stronger assumption (2.5) in the preprint by $(\omega_{1})$.
@@
Let $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing. If $\omega$ satisfies $(\omega_{1})$, then
@@
and then apply [12, Prop. 2.24]. Note that (2.7) is equivalent to the existence of universal constants $C_{1},C,D\geq 1$ such that $C_{1}\omega^{\iota}(s)+D\geq\frac{1}{C_{2}}\min\left\{1,\frac{t^{\beta}}{s^{\beta}}\right\}\omega^{\iota}(s)$ holds for all $s,t\in(0,+\infty)$ (set $C:=\max\{D/C_{1},C_{1}C_{2}\}$).
@@
For the second case we follow the ideas presented by one of the referees in the report: By $(\omega_{1})$, and since $\omega$ is non-decreasing, we get $\omega(2t)\leq L\omega(t)+L$ for some $L>1$ and all $t\geq 0$. By iteration we obtain (see e.g. [28, Sect. 3.3, (3.3.1)])
$\exists\;L>1\;\forall\;n\in\mathbb{N}_{>0}\;\forall\;t\geq 0:\;\;\;\omega(2^{n}t)\leq L^{n}\omega(t)+\sum_{j=1}^{n}L^{j}=L^{n}\omega(t)+\frac{L^{n+1}-L}{L-1}\leq L^{n}\omega(t)+L^{n+1}.$
According to this given $L$ we set $\beta:=\log(L)/\log(2)$, i.e. $2^{\beta}=L$, and so $\beta>0$. Now let $\lambda\geq 1$ be given, choose $n\in\mathbb{N}_{>0}$ such that $2^{n}\leq\lambda<2^{n+1}$ and estimate as follows for all $t\geq 0$:
$\omega(\lambda t)\leq\omega(2^{n+1}t)\leq L^{n+1}\omega(t)+L^{n+2}\leq\lambda^{\beta}(2^{\beta}\omega(t)+4^{\beta}).$
The second estimate is valid since $L^{n+1}=2^{\beta(n+1)}\leq(2\lambda)^{\beta}\Leftrightarrow 2^{n}\leq\lambda$ and $L^{n+2}=2^{\beta(n+2)}\leq(4\lambda)^{\beta}\Leftrightarrow 2^{n}\leq\lambda$. Thus, with $t^{\prime}:=\frac{1}{\lambda t}$ we get
$\forall\;\lambda\geq 1\;\forall\;t^{\prime}>0:\;\;\;\omega^{\iota}(t^{\prime})=\omega(\lambda t)\leq\lambda^{\beta}(2^{\beta}\omega(t)+4^{\beta})=\lambda^{\beta}(2^{\beta}\omega^{\iota}(\lambda t^{\prime})+4^{\beta}).$
Let now $0<t<s$ be arbitrary and with $\lambda=\frac{s}{t}>1$ the above estimate gives
$\left(\frac{s}{t}\right)^{\beta}(2^{\beta}\omega^{\iota}(s)+4^{\beta})=\lambda^{\beta}(2^{\beta}\omega^{\iota}(\lambda t)+4^{\beta})\geq\omega^{\iota}(t),$
and so it follows that
$\forall\;s>t>0:\;\;\;2^{\beta}\omega^{\iota}(s)+4^{\beta}\geq\left(\frac{t}{s}\right)^{\beta}\omega^{\iota}(t)=\min\left\{1,\frac{t^{\beta}}{s^{\beta}}\right\}\omega^{\iota}(t).$
@@
And if there exists some $t_{0}>0$ such that $\omega(t)=0$ for all $t\in[0,t_{0}]$, then $\omega^{\iota}(t)=0$ for all $t\geq t_{0}^{-1}$. Applying the previous estimate to $s:=t_{0}^{-1}$ gives $\omega^{\iota}(\lambda_{1}t+\lambda_{2}t_{0}^{-1})\leq A^{2}\lambda_{1}\omega^{\iota}(t)+A^{2}$ for all $t\in(0,t_{0}]$ and $\lambda_{1},\lambda_{2}\geq 0$ satisfying $\lambda_{1}+\lambda_{2}=1$.
Remark that the proof and the conclusion from [28, Lemma 3.8.1 $(2)$] fails since super-additivity for $\omega^{\iota}$ would imply $2\omega^{\iota}(t)\leq\omega^{\iota}(2t)\leq\omega^{\iota}(t)$ for all $t\in(0,+\infty)$, a contradiction.
@@ 2.6. Comment on weights being defined on $\mathbb{R}^{d}$ or $\mathbb{C}^{d}$
2.7. The growth indices $\gamma(\omega)$ and $\overline{\gamma}(\omega)$
Let in this section $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing and such that $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$; i.e. $\omega$ is a weight function as in [12, Sect. 2.3]. We briefly recall the definition of the growth index $\gamma(\omega)$, see [12, Sect. 2.3] and the references therein. For $\gamma>0$ we say that $\omega$ has property $(P_{\omega,\gamma})$ if
@@ 2.7. The growth indices $\gamma(\omega)$ and $\overline{\gamma}(\omega)$
$\gamma(\omega):=\sup\{\gamma>0:(P_{\omega,\gamma})\;\;\text{is satisfied}\},$
@@
Let $\omega$ be as before. We recall some connections between $\gamma(\omega)$ and sub-additivity-like conditions for $\omega$, see also [29, Sect. 6]:
@@
    $\kappa_{\omega}$ is concave, continuous and we have $\kappa_{\omega}(0)=\omega(0)\geq 0$ (i.e. $\kappa_{\omega}(0)=0$ if and only if $\omega(0)=0$). Hence it is known that also $(\omega_{\operatorname{sub}})$ holds for $\kappa_{\omega}$; see e.g. [28, Lemma 3.8.1 $(1)$].
@@
Analogously, for $\gamma>0$ we say that $\omega$ has property $(\overline{P}_{\omega,\gamma})$ if
(2.11) $\exists\;A>1:\;\;\;\liminf_{t\rightarrow+\infty}\frac{\omega(A^{\gamma}t)}{\omega(t)}>A.$
If $(\overline{P}_{\omega,\gamma})$ holds for some $A>1$, then $(\overline{P}_{\omega,\gamma^{\prime}})$ is satisfied for all $\gamma^{\prime}\geq\gamma$ with the same $A$ since $\omega$ is non-decreasing. Moreover, we can restrict to $\gamma>0$ because for $\gamma\leq 0$ condition $(\overline{P}_{\omega,\gamma})$ is never satisfied for any weight function: $\omega$ is assumed to be non-decreasing and $A>1$. Then set
(2.12) $\overline{\gamma}(\omega):=\inf\{\gamma>0:\;\;(\overline{P}_{\omega,\gamma})\;\;\text{is satisfied}\}.$
We obtain $(0\leq)\gamma(\omega)\leq\overline{\gamma}(\omega)$, see again [12, Sect. 2.3 & 2.4], and via [12, Thm. 2.16 & Cor. 2.17] we get $\overline{\gamma}(\omega)<+\infty$ if and only if $\omega$ satisfies $(\omega_{6})$.
2.8. Comments on different notions of weight functions
In this section we comment in more detail on the different notions of weight functions appearing in the literature; recall the definitions given in the introduction (Section 1).
For the abstract study of weight functions, the assumptions continuity and $\omega(0)=0$ in Definition 1.1 can be removed; e.g. when investigating the growth indices $\gamma(\omega)$ and $\overline{\gamma}(\omega)$; see [12, Sect. 2.2-2.5] and Section 2.7.
Moreover, for a BMT-weight function (see Definition 1.4) we have the following terminology: